Calculate /1.12.13.4ls and Is



R1 = 1kohms→I1



R2 = 1kohms→I2



R3 = 3kohms→I3


I4 R4→ 5K ohms


I5→6Kohms


R6→ 6K ohms


I6 ←


R7 2K ohms

1
Expert's answer
2022-03-28T14:24:18-0400

Assume V=12volt

Parrallel combination


"V_1=V_2=V_3=V_4=V_5=V_6=V_7"

Current

"I_1=\frac{V_1}{R_1}\\I_1=\frac{12}{1\times10^3}=12mA"

"I_2=\frac{V_2}{R_2}\\I_2=\frac{12}{1\times10^3}=12mA"

"I_3=\frac{V_3}{R_3}\\I_3=\frac{12}{3\times10^3}=4mA"

"I_4=\frac{V_4}{R_4}\\I_4=\frac{12}{5\times10^3}=2.4mA"

"I_5=\frac{V_5}{R_5}\\I_5=\frac{12}{6\times10^3}=2mA"

"I_6=\frac{V_6}{R_6}\\I_6=\frac{12}{6\times10^3}=2mA"

"I_7=\frac{V_7}{R_7}\\I_7=\frac{12}{2\times10^3}=6mA"